what is the current when a typical static charge of 0.250 mC moves from your finger to a metal doorknob in 1.00 ms?
1
Expert's answer
2020-02-27T10:09:31-0500

"I=\frac{\Delta q}{\Delta t}=\frac{0.250\cdot 10^{-3}}{1.00\cdot 10^{-3}}=0.25 A"